OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = −5x + 2, solve for f(g(x)) when x = 3....PLEASE HELP

OpenStudy (welshfella):

first, replace the x in g(x)by 3 and work it out:- g(3) = -5(3) + 2 then plug this result into f(x) , 2x + 1

OpenStudy (welshfella):

-5(3) + 2 = -15 + 2 = ?

OpenStudy (anonymous):

-13

OpenStudy (anonymous):

so its -25?

OpenStudy (welshfella):

correct now plug this into 2x + 1 and work it out that's your answer

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

got it thanks!

OpenStudy (welshfella):

yw

hero (hero):

@welshfella what is f(g(x))?

OpenStudy (welshfella):

f(g(x)) = 2(-5x + 2) + 1

hero (hero):

Yes you should explain to the user how to calculate f(g(x)) first before evaluating f(g(3)). The user may need to know how to do that, especially if asked to graph f(g(x)) for example.

hero (hero):

@welshfella

OpenStudy (welshfella):

yes - point taken